OpenStudy (lgbasallote):

a toy rocket engine is securely fastened to a large puck that can glide with negligible friction over a horizontal surface, taken as xy plane. the 4 kg pick has a velocity of 3.00 i m/s at one instant. eight seconds later its velocity is (8i + 10j) m/s. assuming the rocket engine exerts a constant horizontal force find the components of the force and its magnitude

OpenStudy (goformit100):

@cshalvey

OpenStudy (anonymous):

I got F = 5/2 i + 5 j..........I'm not sure if I'm right.

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

we have ------- V(0) = 3i and V(8) = 8i + 10j F = constant ---> F = ma ---> F = mdv/dt ---> Fdt = mdv integration and then we have : F*t = mv + c F*0 = 4*(3i) + c (1) F*8 = 4*(8i + 10j) + c (2) solve for c from equation 1 and put in the equation 2 8F=20i+40j F=5/2i+5j u r right @Vaidehi09

OpenStudy (lgbasallote):

uhmm what happened here? will someone be kind enough to explain the steps? haha sorry...like i said im new to this

OpenStudy (foolaroundmath):

Given: acceleration = constant. Let it be \(a = x\hat{i}+y\hat{j}\) We know that for constant acceleration, \(\vec{v} = \vec{u} + \vec{a}t\) where \(\vec{v}\) is the final velocity, \(\vec{u}\) is the initial velocity and t = time taken. Given: \(\vec{u} = 3\hat{i}, \vec{v} = 8\hat{i}+10\hat{j}, t = 8\) Substituting, we get: \(\vec{a} = (5\hat{i}+10\hat{j})/8\) Now, force \(\vec{F} = m\vec{a}\) and m = 4kg So, \(\vec{F} = (5/2)\hat{i} + 5\hat{j}\) \(|\vec{F}| = \sqrt{(5/2)^{2}+5^{2}}= 5\sqrt{5}/2\)